# Test-Time Modification: Inverse Domain Transformation for Robust Perception

### v2

@@ Abstract
Generative foundation models contain broad visual knowledge and can produce diverse image variations, making them particularly promising for advancing domain generalization tasks. While they can be used for training data augmentation, synthesizing comprehensive target-domain variations remains slow, expensive, and incomplete. We propose an alternative: using diffusion models at test time to map target images back to the source distribution where the downstream model was trained. This approach requires only a source domain description, preserves the task model, and eliminates large-scale synthetic data generation. We demonstrate consistent improvements across segmentation, detection, and classification tasks under challenging environmental shifts in real-to-real domain generalization scenarios with unknown target distributions. Our analysis spans multiple generative and downstream models, including an ensemble variant for enhanced robustness. The method achieves substantial relative gains: 137% on BDD100K-Night, 68% on ImageNet-R, and 62% on DarkZurich.
[Uncaptioned image][Uncaptioned image]
@@ 1 Introduction
Despite major advances in vision architectures and training strategies, model performance still relies strongly on the quality and coverage of the training data. Predictions tend to be reliable when test images follow the same distribution as the training set, but even moderate distribution shifts can lead to a clear drop in accuracy. To tackle this, the field of domain generalization aims to train models that remain robust to unseen target domains without assuming access to data or annotations from those domains. Most existing approaches fall into two categories (Fig. 2b,c).
@@
The first one “Unspecific Image Augmentation” modifies the training data in broad, unspecific ways: typically through heavy augmentations such as blur, noise, or color jitter, with the hope that the model will generalize to new domains. The second approach, “Generative Training Image Augmentation” leverages recent diffusion-based generative models to synthesize additional training data. Using text prompts, these methods attempt to approximate potential target domains and generate corresponding images. However, estimating all relevant target domains in advance is difficult, and the diversity of synthesized data remains limited.
@@
In this work, we introduce a third direction. Instead of expanding the training distribution, we use modern generative models to compute an inverse transformation at test time. The goal is to map a target-domain image back toward the source-domain distribution on which the discriminative model was trained. This idea is related to test-time normalization techniques such as adjusting BatchNorm statistics, but diffusion-based transformations allow for far richer and more expressive changes. Because the discriminative model performs best on its original training distribution, this inverse mapping improves prediction quality without modifying the model itself. Importantly, computing this inverse transformation requires only a description of the source domain, which is readily available. In contrast, expanding the training distribution with synthetic data would require anticipating all relevant target domains, which is a far harder problem.
We emphasize that our approach is not intended as a runtime optimization for deployment scenarios, as it requires inference through large generative models that introduce significant computational overhead. Rather, we present this work as a proof of principle that demonstrates how the vast world knowledge encoded in foundation models can be effectively harnessed to improve traditional computer vision tasks. Our method showcases a paradigm where generative models serve as powerful domain translators, unlocking their learned representations to bridge distribution gaps without any task-specific training. This demonstration opens new research directions for understanding and utilizing the complementary strengths of generative and discriminative models, even if practical deployment would require further innovations in efficiency.
@@
We evaluate our inverse domain transformation on domain generalization benchmarks for semantic segmentation, detection, and classification. Our test-time modification (TTM) pipeline improves performance across all settings without finetuning either the discriminative or generative models, offering a simple, task-agnostic approach to domain generalization. Our main contributions are as follows:
@@
  • Quantitative evaluation and analysis of Test-Time Modification (TTM) as a new paradigm that requires no retraining of generative or discriminative models
  • State-of-the-art results for various pretrained models on benchmarks such as BDD [yu2020bdd100k], DarkZurich [sakaridis2019guided], ACDC [sakaridis2021acdc], ImageNet-R [hendrycks2021many]
@@ 2 Related Work
Constraining the distributions. Early methods constrain input distributions via normalization in latent space, using batch-norm [SEGU2023109115] or instance normalization [Pan2018IBNNet] with test-time parameters from target images. Subsequent works introduce whitening [pan2019Whitening] to decorrelate channel representations and reduce style information, with various improvements for guided whitening [Choi2021, xu2022dirl, Peng2022semanticaware]. Our approach instead leverages generative models’ world knowledge to map target distributions back to source, providing more robust transformations that handle aleatoric uncertainty.
@@
Data augmentation and randomization. Extending source distributions to include target domain information has been widely explored, from basic augmentations (Gaussian blur, color jitter) to domain randomization [Yue2019DomainRandomization] using random style transformations. Several works [peng2021DomRandGlobal, kim2021DomRandwedge] leverage auxiliary domains (paintings, web images) for style randomization, while Huang et al. [Huang2021DomRand] perform randomization in frequency space. Zhong et al. [Zhong2022advStyle] proposed an adversarial style augmentation and Zhao et al. [Zhao2022Halluc] a style hallucination module for extending the source domain.
@@
Prior knowledge for source domain extension. Generative models like Stable Diffusion [rombach2022highresolutionimagesynthesislatent] enable synthetic target domain creation using prior knowledge. Niemeijer et al.[Niemeijer2024WACV] generate pseudo-target domains with weather and traffic variations for autonomous driving, then apply unsupervised domain adaptation [Schwonberg2023Survey]. Benigmim et al. [Benigmim_2024_CVPR] utilize the prior knowledge contained in LLMs to create meaningful prompts the generation of relevant synthetic data. ControlNet variants [zhang2023adding, li2024adversarial, Niemeijer2025ICCV, Li2024ControlNetIC] enable supervised training on synthetic data by conditioning on both labels and target domain descriptions. However, limited prior knowledge of prevents comprehensive coverage of all relevant domains that might be encountered at test-time. In contrast, we argue that using these generative models is more efficient when done in reverse: we describe the source domain and transform target images back to the source distribution.
@@
Inverse transformations. Prior inverse transformation methods use style transfer via Fourier transforms [Termöhlen2021Continual] or GANs [mütze2022Iverse], but have required target domain access to determine style parameters. The approach of Yu et al. [Yu_2023_CVPR] is closest to ours. They proposed an inverse transformation to the source domain based on a source domain trained unconditional diffusion model. We instead propose a new paradigm that leverages world knowledge from generative models like Flux1. Kontext [labs2025flux1kontextflowmatching] and Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport] to perform arbitrary transformations using only text descriptions. Beyond style transfer, these models can interpret image content and even reduce aleatoric uncertainty, as we describe in Sec.3.1.
@@ 3 Methods
Generative models like Stable Diffusion [rombach2022highresolutionimagesynthesislatent, wu2023datasetdm] are trained on large-scale web data containing diverse image distributions, enabling them to capture extensive world knowledge. Recent works [Niemeijer2024WACV, kupyn2024dataset, wu2023diffumask] have exploited this capability to augment training data distributions. In these approaches, a diffusion generative model $G$ is guided by text prompt $t_{k}$ to create synthetic images $x^{PT}_{n}$ that approximate target domain characteristics:
@@
Test-Time Image Modification. Instead of generating pseudo-target domains during training, we propose using generative models for image editing (e.g., Flux.1 Kontext [labs2025flux1kontextflowmatching] or Qwen-Image-Edit [wu2025qwenimagetechnicalreport]) to transform target domain images back to the source domain at test time:
$x_{n}^{PS}=G(x_{n}^{T},t^{S})$ (2)
where $t^{S}$ describes the source domain distribution.
@@
Since $f_{\theta}$ was trained on the source domain distribution, predictions on $x_{n}^{PS}$ should be more accurate than those obtained directly from $x_{n}^{T}$. Crucially, this approach requires no additional training or fine-tuning of the discriminative model.
@@
Figure 3: Overview of our test-time modification pipeline via inverse domain transfer. The process begins with formulating a meta-prompt that describes the desired transformation. An MLLM processes this meta-prompt to generate an effective prompt $t^{S}$ for the I2I generation model. During test time, the I2I model transforms target domain images to the source domain. Finally, the pre-trained discriminative model processes both images, with task-specific fusion strategies applied as needed.
Generating Effective Source Domain Prompts. To obtain an effective source domain prompt $t^{S}$, we employ the two-stage pipeline illustrated in Figure 3.
Step 1: Meta-Prompt Formulation. We construct a comprehensive meta-prompt: a human-written text, denoted as $t_{\text{meta}}$, designed to generate another prompt. This meta-prompt instructs a multimodal large language model (MLLM), such as GPT-5 [openai_gpt5], to generate a prompt suitable for inverse domain transformation using an Image-to-Image (I2I) generation model. The meta-prompt $t_{\text{meta}}$ integrates the following components:
@@
  • Domain context: Application-specific information (e.g., “road driving scenes”)
@@ 3.1 Intuition: Reducing Aleatoric Uncertainty
This allows our test-time modification to mitigate effects such as snow cover, raindrops, or low-light conditions (see Figures 4 and 7). Importantly, this reduction in aleatoric uncertainty would not be achievable by synthesizing pseudo-target domains during training, as aleatoric uncertainty is fundamentally tied to the input rather than the model.
@@ 4 Experiments
Table 1: Semantic segmentation performance (mIoU in %) on different domain generalization benchmarks with and without (base model) using image to image generation model based test-time modification (TTM). CS refers to Cityscapes and Geo. shift refers to geographic shift.
Method DeepLabV3+ [ResNet-101-d8] OCRNet [HRNetV2-W48] Segformer [MiT-B1] Segformer [MiT-B5] Mask2Former [Swin-L] Avg mIoU $\Delta\uparrow$ (%)
@@
Across all downstream tasks, we use a total of six target datasets for evaluation and two datasets for training our source domain models. For semantic segmentation, we use the clear weather Cityscapes [cordts2016cityscapes] as our source dataset. Whereas we evaluate on adverse weather and nighttime datasets - ACDC [sakaridis2021acdc], BDD100K [yu2020bdd100k], BDD100K-Night [sakaridis2020map, yu2020bdd100k], and Dark Zurich [sakaridis2019guided]. For the object detection task, we again use Cityscapes to train our source model and evaluate on a randomly sampled subset of 501 BDD100K nighttime images with object detection annotations. We refer to this validation set as BDD100K-Night-Det. For image classification, we use the ImageNet-1K [deng2009imagenet] dataset as our source domain and ImageNet-R [hendrycks2021many] dataset as our target set. The latter contains difficult, rendered image versions of a subset of 200 categories. Refer to the appendix for more information on the various datasets.
@@
For our semantic segmentation experiments, we utilize both convolutional neural networks (CNNs) and transformer-based models of varying sizes. CNN based models include DeepLabV3+ [chen2018encoder] and OCRNet [yuan2020object] whereas transformer based models include Segformer [xie2021segformer] and Mask2Former [cheng2022masked]. All these models are pre-trained on the Cityscapes dataset and are directly evaluated on the validation set of the target dataset. For object detection evaluation, we use the Faster-RCNN [ren2015faster] and Mask-RCNN [he2017mask] models. For the image classification task, we use the ResNet-50 and ResNet-152 models [he2016deep]. For image generation, we utilize two different image-to-image (I2I) generation models as part of TTM, including a commercial model, Flux.1 Kontext Max and an open-source model Qwen-Image-Edit-2509 [QIE-2509]. We also do an ablation study on additional I2I generation models. More details on different models are provided in the appendix.
@@ 4.1 Semantic Segmentation
Image Mask2Former w/o TTM Test-Time Modified Image Mask2Former w/ TTM Ground Truth
road sidew. build. wall fence pole tr. light tr. sign veget. terrain sky person rider car truck bus train m.bike bike n/a.
@@
(b) Handcrafted Prompting
(c) Metaprompting GPT-5
(d) MP w/ TTM Formulation
Figure 6: Comparing prompting strategies for inverse domain transfer using QIE-2509. MP in (d) refers to Metaprompting GPT-5
@@
Figure 7: Qualitative performance comparison for domain generalized object detection on Cityscapes $\rightarrow$ BDD100KNight-Det using Faster-RCNN - without and with using TTM. Three different examples from BDD100KNight-Det dataset. (Zoom recommended)
@@
w/ TTM 43.7 $\pm$ 1.2 43.1 $\pm$ 1.3 42.9 $\pm$ 1.6 47.5 $\pm$ 1.3 50.5 $\pm$ 1.2
We evaluate on the domain generalized semantic segmentation task in the autonomous driving context, focusing on weather and daytime shifts, as they cause the most significant domain gap in autonomous driving applications.
@@
Table 1 shows a performance comparison of our Test-Time Modification (TTM) method on the domain generalized semantic segmentation task. As shown in the table, using TTM yields significant improvements in mIoU across various datasets and segmentation models. This performance jump is even more significant with smaller models, such as DeepLabV3+, OCRNet, and Segformer [MiT-B1]. This demonstrates that weaker models can be made more robust by using TTM. An interesting observation that supports this reasoning is that the weakest DeepLabV3+ model, when used with TTM, outperforms a much stronger Segformer [MiT-B5] model for the CS $\rightarrow$ ACDC benchmark. The same trend persists with CS $\rightarrow$ DarkZurich and CS $\rightarrow$ BDD100K-Night. Also, for CS $\rightarrow$ BDD100K, Flux1. Kontext Max outperforms QIE-2509 likely because it is more robust to geographical shifts, which is the primary change in BDD100K w.r.t Cityscapes. Qualitative results for domain generalized semantic segmentation on different target datasets are shown in Fig. 4.
Ablation
@@
We perform an ablation study with QIE-2509 in Table 2, where we generate 10 different Dark Zurich validation sets (for 10 random seeds with the same prompt) using TTM and report the average mIoU across these generations with standard deviation. We try to study the scale of randomization in these I2I models. Considering a small validation set (50 images), the deviation is reasonable, and we are still able to make significant improvements. We perform another ablation in Table 3 where we use five different I2I generation models and compare the segmentation performance on the BDD100K-Night dataset. We use the same prompt for all generative models, obtained during the TTM process. Flux. 1 Kontext Max performs the best when used for TTA. We show a comparison of the generation quality of test-modified images between different generation models in Fig. 5. As can be seen, all models are capable of performing night-to-day transfer with good fidelity. However, Flux.1 Kontext Max maintains the best semantic consistency among all models, as highlighted in the figure. In another ablation study, we show the effect of correct prompting in Fig. 6 by generating images using QIE-2509. Handcrafted short prompts, such as ”remove snow from the image,” often do not yield the desired output. Metaprompting an MLLM, such as GPT-5, is a better strategy; however, without a precise query, an MLLM can generate weak prompts. As we see in Fig. 6(c), the car on the road also gets removed, resulting in a significant layout mismatch. In contrast, our query formulation from the TTM process, as described in Section 3, results in complete snow removal while preserving the semantic layout.
@@
Table 3: CS → BDD100K-Night: segmentation performance comparison for different I2I generative models with Mask2Former segmentation model.
w/o Flux1. Nano- SeedEdit QIE- Flux1.
TTM Kont Dev [labs2025flux1kontextflowmatching] Banana [comanici2025gemini] 3.0 [wang2025seededit] 2509 Kont Max
mIoU 40.6 42.3 43.5 45.3 47.1 49.1
@@
person rider car truck bus motor. bicycle mAP@50 $\Delta\uparrow$ (%)

### v4

@@ Abstract
Generative foundation models contain broad visual knowledge and can produce diverse image variations, making them particularly promising for advancing domain generalization tasks. They can be used for training data augmentation, but synthesizing comprehensive target-domain variations remains slow, expensive, and incomplete. We propose an alternative: using diffusion models at test time to map target images back to the source distribution where the downstream model was trained. This approach requires only a source domain description, preserves the task model, and eliminates large-scale synthetic data generation. We demonstrate consistent improvements across segmentation, detection, and classification tasks under challenging environmental shifts in real-to-real domain generalization scenarios with unknown target distributions. Our analysis spans multiple generative and downstream models, including an ensemble variant for enhanced robustness. The method improves BDD100K-Night-Det mAP@50 from 10.2 to 31.8, ImageNet-R top-1 from 36.1 to 60.8, and DarkZurich mIoU from 28.6 to 46.3.
@@ 1 Introduction
Despite major advances in vision architectures and training strategies, model performance still relies strongly on the quality and coverage of the training data. Predictions tend to be reliable when test images follow the same distribution as the training set, but even moderate distribution shifts can lead to a clear drop in accuracy. To tackle this, the field of domain generalization aims to train models that remain robust to unseen target domains without assuming access to data or annotations from those domains. Most existing approaches fall into two categories (Fig.˜2b,c).
@@
The first one “Unspecific Image Augmentation” modifies the training data in broad, unspecific ways: typically through heavy augmentations such as blur, noise, or color jitter, with the hope that the model will generalize to new domains. The second approach, “Generative Training Image Augmentation” leverages recent diffusion-based generative models to synthesize additional training data. Using text prompts, these methods attempt to approximate potential target domains and generate corresponding images. Estimating all relevant target domains is difficult, and the diversity of synthesized data remains limited.
@@
In this work, we introduce a third direction. Instead of expanding the training distribution, we use modern generative models to compute an inverse transformation at test time. The goal is to map a target-domain image back towards the source-domain distribution on which the discriminative model was trained. This idea is related to test-time normalization techniques such as adjusting BatchNorm statistics, but diffusion-based transformations allow for far richer and more expressive changes. Unlike test-time adaptation (TTA), our approach does not require target-domain statistics. Because the discriminative model performs best on its original training distribution, this inverse mapping improves prediction quality without modifying the model itself. Importantly, computing this inverse transformation requires only a description of the source domain, which is readily available. In contrast, expanding the training distribution with synthetic data requires anticipating all relevant target domains, which is a far harder problem. Moreover, compared to retraining-based domain adaptation pipelines, our method trades expensive offline retraining for a lightweight inference-time transformation.
Our approach introduces a single image-to-image transformation step at inference time, implemented using modern generative models. Recent advances in efficient architectures and optimized inference enable this step to run in near real time while preserving accuracy (Sec.˜4.3.1).Because the transformation is independent of the downstream model, improvements in generative model efficiency directly translate to faster deployment. Our method showcases a new paradigm where generative models serve as powerful domain translators, unlocking their learned representations to bridge distribution gaps without any task-specific training. This opens new research directions for understanding and utilizing the complementary strengths of generative and discriminative models.
@@
We evaluate our inverse domain transformation on domain generalization benchmarks for semantic segmentation, detection, and classification. Our test-time modification (TTM) pipeline improves performance across all settings without finetuning either the discriminative or generative models, offering a simple, task-agnostic approach to domain generalization. Our main contributions are:
@@
  • Quantitative evaluation and analysis of Test-Time Modification (TTM) as a new paradigm that requires no retraining or fine-tuning
  • State-of-the-art results for various pretrained models on benchmarks such as BDD [43], DarkZurich [30], ACDC [32], ImageNet-R [12]
@@ 2 Related Work
Constraining the distributions. Early methods constrain input distributions via normalization in latent space, using batch-norm [34] or instance normalization [24] with test-time parameters from target images. Subsequent works introduce whitening [25] to decorrelate channel representations and reduce style information, with various improvements for guided whitening [6, 42, 26]. Our approach instead leverages generative models’ world knowledge to map target distributions back to source, providing more robust transformations that handle aleatoric uncertainty.
@@
Data augmentation and randomization. Extending source distributions to include target domain information has been widely explored, from basic augmentations (Gaussian blur, color jitter) to domain randomization [46] using random style transformations. Several works [27, 14] leverage auxiliary domains (paintings, web images) for style randomization, while Huang et al. [13] perform randomization in frequency space. Zhong et al. [49] proposed an adversarial style augmentation and Zhao et al. [48] a style hallucination module.
@@
Prior knowledge for source domain extension. Generative models like Stable Diffusion [29] enable synthetic target domain creation using prior knowledge. Niemeijer et al.[22] generate pseudo-target domains with weather and traffic variations for autonomous driving, then apply unsupervised domain adaptation [33]. Benigmim et al. [2] utilize the prior knowledge contained in LLMs to create meaningful prompts the generation of relevant synthetic data. ControlNet variants [47, 18, 21, 17] enable supervised training on synthetic data by conditioning on both labels and target domain descriptions. However, limited prior knowledge of prevents comprehensive coverage of all relevant domains that might be encountered at test-time. In contrast, we argue that using these generative models is more efficient when done in reverse: we describe the source domain and transform target images back to the source distribution.
@@
Inverse transformations. Prior inverse transformation methods use style transfer via Fourier transforms [35] or GANs [20], but require target domain access to determine style parameters. Our transformation does not require additional information from the given target domain image, distinguishing it from test time adaptation (TTA). We therefore do not compare to TTA approaches in the experiments. Further, while generalist video models can solve perception tasks zero-shot [37], they are computationally expensive and still underperform task-specific discriminative models that TTM is designed to enhance — by employing generative models as a targeted pre-processor that bridges domain gap at test time with reliable single-pass inference. The approach of Yu et al. [44] is closest to ours. They proposed an inverse transformation to the source domain based on a source domain trained unconditional diffusion model. However, this strategy relies on a source-trained diffusion model and is studied primarily in a classification-only setting. In contrast, we introduce a new paradigm that leverages the world knowledge encoded in off-the-shelf foundation editing models (e.g., Flux1. Kontext and Qwen-Image-Edit-2509) together with a text-defined source description, yielding a domain-agnostic, plug-and-play inverse transformation that generalizes across segmentation, detection, and classification. Kontext [16] and Qwen-Image-Edit-2509 [38] to perform arbitrary transformations using only text descriptions. Beyond style transfer, these models can interpret image content and even reduce aleatoric uncertainty, as we describe in Sec.˜3.1.
@@ 3 Methods
Generative models like Stable Diffusion [29, 39] are trained on large-scale web data containing diverse image distributions, enabling them to capture extensive world knowledge. Recent works [22, 15, 40] have exploited this capability to augment training data distributions. In these approaches, a diffusion generative model $G$ is guided by text prompt $t_{k}$ to create synthetic images $x^{PT}_{n}$ that approximate target domain characteristics:
@@
Definition: Test-Time Modification (TTM) Instead of generating pseudo-target domains during training, we propose using generative models for image editing (e.g., Flux.1 Kontext [16] or Qwen-Image-Edit [38]) to transform target domain images back to the source domain at test time: $x_{n}^{PS}=G(x_{n}^{T},t^{S})$ (2) where $t^{S}$ describes the source domain distribution. TTM assumes no access to target-domain data beyond the current test image and performs no model adaptation.
@@
Since $f_{\theta}$ was trained on the source domain distribution, predictions on $x_{n}^{PS}$ should be more accurate than those obtained directly from $x_{n}^{T}$. Crucially, this approach requires no additional training of the discriminative model.
@@
Figure 3: Overview of our test-time modification pipeline doing inverse domain transfer. The process begins with formulating a meta-prompt that describes the desired transformation. An MLLM processes this meta-prompt to generate an effective prompt $t^{S}$ for the I2I generation model. During test time, the I2I model transforms target domain images to the source domain. Finally, the pre-trained discriminative model processes both images, with task-specific fusion strategies applied as needed.
Generating Effective Source Domain Prompts. To obtain an effective source domain prompt $t^{S}$, we employ the two-stage pipeline illustrated in Fig.˜3.
Step 1: Meta-Prompt Formulation. We construct a comprehensive meta-prompt: a human-written text, denoted as $t_{\text{meta}}$, designed to generate another prompt. This meta-prompt instructs a multimodal large language model (MLLM), such as GPT-5 [23], to generate a prompt suitable for inverse domain transformation using an Image-to-Image (I2I) generation model. The meta-prompt $t_{\text{meta}}$ integrates the following components:
@@
  • Domain context: Application-specific information (e.g. “road driving scenes”)
@@ 3.1 Intuition: Reducing Aleatoric Uncertainty
This allows our test-time modification to mitigate effects such as snow cover, raindrops, or low-light conditions (see Fig.˜4 and Fig.˜7). Importantly, this reduction in aleatoric uncertainty would not be achievable by synthesizing pseudo-target domains during training, as aleatoric uncertainty is fundamentally tied to the input rather than the model.
@@ 4 Experiments
Table 1: Semantic segmentation performance (mIoU in %) on different domain generalization benchmarks with and without (base model) using test-time modification (TTM). CS refers to Cityscapes and Geo. shift refers to geographic shift.
[…]
@@
Across all downstream tasks, we use a total of six target datasets for evaluation and two datasets for training our source domain models. For semantic segmentation, we use the clear weather Cityscapes [8] as our source dataset. Whereas we evaluate on adverse weather and nighttime datasets - ACDC [32], BDD100K [43], BDD100K-Night [31, 43], and Dark Zurich [30]. For the object detection task, we again use Cityscapes to train our source model and evaluate on a randomly sampled subset of 501 BDD100K nighttime images with object detection annotations. We refer to this validation set as BDD100K-Night-Det. For image classification, we use the ImageNet-1K [9] dataset as our source domain and ImageNet-R [12] dataset as our target set. The latter contains difficult, rendered image versions of a subset of 200 categories. Refer to the appendix for more information on the various datasets.
@@
For our semantic segmentation experiments, we utilize both convolutional neural networks (CNNs) and transformer-based models of varying sizes. CNN based models include DeepLabV3+ [4] and OCRNet [45] whereas transformer based models include Segformer [41] and Mask2Former [5]. All these models are pre-trained on the Cityscapes dataset and are directly evaluated on the validation set of the target dataset. For object detection evaluation, we use the Faster-RCNN [28] and Mask-RCNN [10] models. For the image classification task, we use the ResNet-50 and ResNet-152 models [11]. For image generation, we mainly utilize two different image-to-image (I2I) generation models as part of TTM, including a commercial model, Flux.1 Kontext Max and an open-source model Qwen-Image-Edit-2509 [QIE-2509]. We also do an ablation study on additional I2I generation models including Flux.2 variants [3]. More details on different models are provided in the appendix.
@@ 4.1 Semantic Segmentation
Original Image w/o TTM TTM Image w/ TTM Ground Truth
[…]
@@
(b) Handcrafted
(c) Metaprompt LLM
(d) Ours (TTM)
Figure 6: Comparing prompting strategies for inverse domain transfer using QIE-2509. Handcrafted prompts (a) often fail to produce the desired inverse transformation. Metaprompting an MLLM like GPT-5 (c) comes closer to achieving the desired inverse transformation but still deviates while often compromising the semantic consistency, e.g. missing a car on the road. Metaprompting an MLLM with our TTM formulation (d) carries concrete guidance necessary to achieve the desired transformation without trading off the semantic consistency.
@@
Figure 7: Qualitative performance comparison for domain generalized object detection on Cityscapes $\rightarrow$ BDD100KNight-Det using Faster-RCNN - without and with using TTM. Three different examples from BDD100KNight-Det dataset.
@@
w/ TTM 43.7 $\pm$ 1.3 43.1 $\pm$ 1.3 42.9 $\pm$ 1.6 47.5 $\pm$ 1.3 50.5 $\pm$ 1.2
@@
We evaluate on the domain generalized semantic segmentation task in the autonomous driving context, focusing on weather and daytime shifts, as they cause the most significant domain gap in autonomous driving applications. Table˜1 shows a performance comparison of our Test-Time Modification (TTM) method on the domain generalized semantic segmentation task. As shown in the table, using TTM yields significant improvements in mIoU across various datasets and segmentation models. This performance jump is even more significant with smaller models, such as DeepLabV3+, OCRNet, and Segformer [MiT-B1]. This demonstrates that weaker models can be made more robust by using TTM. An interesting observation that supports this reasoning is that the weakest DeepLabV3+ model, when used with TTM, outperforms a much stronger Segformer [MiT-B5] model for the CS $\rightarrow$ ACDC benchmark. The same trend persists with CS $\rightarrow$ DarkZurich and CS $\rightarrow$ BDD100K-Night. Also, for CS $\rightarrow$ BDD100K, Flux1. Kontext Max outperforms QIE-2509 likely because it is more robust to geographical shifts, which is the primary change in BDD100K w.r.t Cityscapes. Qualitative results for domain generalized semantic segmentation on different target datasets are shown in Fig.˜4.
4.1.1 Ablations.
@@
We perform an ablation study with QIE-2509 in Tab.˜2, where we generate 10 different Dark Zurich validation sets (for 10 random seeds with the same prompt) using TTM and report the average mIoU across these generations with standard deviation. We try to study the scale of randomization in these I2I (Image-to-Image) generation models. Considering a small validation set (50 images), the deviation is reasonable, and we are still able to make significant improvements. We perform another ablation in Tab.˜3 where we use five different I2I generation models and compare the segmentation performance on the BDD100K-Night dataset. We use the same prompt for all generative models, obtained during the TTM process. Flux. 1 Kontext Max performs the best when used for TTM. We show a comparison of the generation quality of test-modified images between different generation models in Fig.˜5. As can be seen, all models are capable of performing night-to-day transfer with good fidelity. However, Flux.1 Kontext Max maintains the best semantic consistency among all models, as highlighted in the figure. In another ablation study, we show the effect of correct prompting in Fig.˜6 by generating images using QIE-2509. Handcrafted short prompts, such as "remove snow from the image," often do not yield the desired output. Metaprompting an MLLM, such as GPT-5, is a better strategy; however, without a precise query, an MLLM can generate weak prompts. As we see in Fig.˜6(c), the car on the road also gets removed, resulting in a significant layout mismatch. In contrast, our query formulation from the TTM process, as described in Fig.˜3, results in complete snow removal while preserving the semantic layout. We also performed an ablation to show the effectiveness of the probability mean fusion step described in Sec.˜3. Table˜4 shows the ablation results for semantic segmentation. Across all evaluated domain shifts, fusion consistently improves TTM performance. Segmentation allows simple pixel-level fusion; whereas detection fusion requires box matching to avoid duplicates and adds task-specific complexity. Conversely, image classification lacks the detailed spatial constraints that necessitate fusion step in the first place. Thus, we only apply the fusion step for semantic segmentation task.
@@
Table 3: CS → BDD100K-Night: segmentation performance comparison for different I2I generative models with Mask2Former segmentation model. Abbreviations: F1.K.Dev=FLUX.1 Kontext [Dev]; NB=Nano-Banana; SE3=SeedEdit 3.0; QIE=Qwen-Image-Edit-2509; F2=FLUX.2 [Dev]; F1.K.Max=FLUX1. Kontext [Max].
w/o TTM F1.K.Dev [16] NB [7] SE3 [36] QIE [38] F2 [3] F1.K.Max [16]
mIoU 40.6 42.3 43.5 45.3 47.1 48.7 49.1
Table 4: Fusion step ablation (mIoU) across target domains using Cityscapes-pretrained Mask2Former using Flux1. Kontext Max. Base model implies no TTM.
Method ACDC Dark Zurich BDDNight BDD
Base Model 60.5 40.6 40.6 58.0
w/o Fusion 60.6 48.7 46.7 55.7
w/ Fusion 64.3 51.7 49.1 58.7
@@
person rider car truck bus motor. bicycle mAP@50 $\Delta\uparrow$